Write the equation of a line that passes through the 2 points (3,4) and (1,-2)
3,4 2/6= m=1/3 y-4=1/3(x-3)
1,-2 y-4=1/3x-1
+4 +4

y=1/3x+3 Is this the correct answer???

Answer by mananth(16949)

x1 y1 x2 y2
3 4 1 -2

slope m = (y2-y1)/(x2-x1)
( -2 - 4 )/( 1 - 3
( -6 / -2 )
m= 3
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point ( 3 , 4 ) in
Y = m x + b
4.00 = 3 * 3 + b
4.00 = 9 + b
b= 4 + -5
b= 9
So the equation will be
Y = 3 x -5
m.ananth@hotmail.ca